The function C(F)= 5/9(F-32) converts a temperature in degrees Fahrenheit,F, to a temperature in degrees Celsius,C. The function K(C)=C+273, converts a temperature in degrees Celsius to a temperature in fegrees kelvins,K.
Find a function that convers a temperature in degrees Fahrenheit to a temperature in Kelvins.

Answer by CharlesG2(834)

The function C(F)= 5/9(F-32) converts a temperature in degrees Fahrenheit,F, to a temperature in degrees Celsius,C. The function K(C)=C+273, converts a temperature in degrees Celsius to a temperature in fegrees kelvins,K.
Find a function that convers a temperature in degrees Fahrenheit to a temperature in Kelvins
C(F) = 5/9 * (F-32) this gives Celsius from Fahrenheit
K(F) = C(F) + 273 = 5/9 * (F-32) + 273 this gives Kelvin from Fahrenheit